OpenStudy (anonymous):

If a town's population data was interpreted with this equation for the line of fit, P = 519Y + 23,100, what is the population of the town five years later after the original date of the data

OpenStudy (anonymous):

Assuming \(Y\) represents time, when \(Y=0\) you have your initial population: \[\text{initial pop.}=519(0)+23,100=23,100\] Also assuming time is given in units of years, the population of the town five years later would occur when \(Y=5\): \[\text{pop. after 5 years}=519(5)+23,100=\cdots\]